OpenStudy (anonymous):

Pleeassse! A graph shows a production of coal in thousands from a scale of 0 -105 covering a 6 month period. Trend A shows a total standard and non standard quality and Trend B shows a standard production of coal in in any month. Question in July the Total standard and non standard was 85000 and Standard 65000. how would i work out the percentage of the non standard production as a percentage?

OpenStudy (anonymous):

Standard/Total = 65000/85000 = 65/80 = 13/16 = 0.8125, or 81.25%